# Refining the WASP-132 multi-planetary system: discovery of a cold giant planet and mass measurement of a hot super-Earth

### v1

@@
Refining the WASP-132 multi-planetary system: discovery of a cold giant planet and mass measurement of a hot super-Earth
Hot Jupiters generally do not have nearby planet companions, as they may have cleared out other planets during their inward migration from more distant orbits. This gives evidence that hot Jupiters more often migrate inward via high-eccentricity migration due to dynamical interactions between planets rather than more dynamically cool migration mechanisms through the protoplanetary disk. Here we further refine the unique system of WASP-132 by characterizing the mass of the recently validated 1.0-day period super-Earth WASP-132 c (TOI-822.02) interior to the 7.1-day period hot Jupiter WASP-132 b. Additionally, we announce the discovery of a giant planet at a 5-year period (2.7 AU). We also detect a long-term trend in the radial velocity data indicative of another outer companion. Using over nine years of CORALIE RVs and over two months of highly-sampled HARPS RVs, we determine the masses of the planets from smallest to largest orbital period to be Mc = , Mb = , and M = , respectively. Using TESS and CHEOPS photometry data we measure the radii of the two inner transiting planets to be and . WASP-132 is a unique multi-planetary system in that both an inner rocky planet and an outer giant planet are in a system with a hot Jupiter. This suggests it migrated via a more rare dynamically cool mechanism and helps to further our understanding of how hot Jupiter systems may form and evolve.
@@
Some previous studies found that most hot Jupiters do not have nearby planet companions. This was based on nondetections of additional planets in radial velocity (RV) data (e.g., Wright et al., 2009) and photometry data by looking for both additional transit signals (e.g., Steffen et al., 2012; Huang et al., 2016; Hord et al., 2021) and transit timing variations (e.g., Steffen et al., 2012; Wang et al., 2021; Ivshina & Winn, 2022). The lack of detected planets may be a result of highly eccentric giant planets clearing out any low-mass inner planets in the system during their migration inward (Mustill et al., 2015). These results suggest that high-eccentricity migration due to dynamical interactions between planets (e.g., Rasio & Ford, 1996; Weidenschilling & Marzari, 1996; Lin & Ida, 1997) may be a more common migration mechanism for gas giant planets compared to disk-driven migration (e.g., Goldreich & Tremaine, 1980; Ward, 1997; Baruteau et al., 2014).
@@
However, Wu et al. (2023) recently searched for transit timing variations across the full four-year Kepler (Borucki et al., 2010) data set and found that at least 126% of hot Jupiters have a nearby companion. Continued surveys and recent advancements in the precision of both photometry and RV observations have allowed the detection of some of these companions including WASP-47 (Becker et al., 2015), Kepler-730 (Zhu et al., 2018; Cañas et al., 2019), TOI-1130 (Huang et al., 2020), WASP-148 (Hébrard et al., 2020), WASP-132 (Hord et al., 2022), and TOI-2000 (Sha et al., 2022). These companions give strong evidence that these hot Jupiters migrated via quiescent mechanisms that are dynamically cool and allow nearby planetary companions to remain in the system.
@@
Giant long-period planets have also been detected in systems with small close-in planets (e.g., Santos et al., 2016). Several studies have worked to determine the occurrence of long-period ( 1 year) giant (Mp 0.3 MJup/95 M⊕; Rp 0.45 RJup/5 R⊕) planets, also known as cold Jupiters, in systems that contain close-in (100 days) small (Rp 4 R⊕; Mp 30 M⊕) planets. These studies have varying results with some suggesting a positive correlation between close-in small planets and cold Jupiters (e.g., Zhu & Wu, 2018; Bryan et al., 2019; Herman et al., 2019; Rosenthal et al., 2022) while others find a negative correlation (e.g., Barbato et al., 2018; Bonomo et al., 2023). A positive correlation suggests that these two planet populations do not directly compete for solid material (e.g., Zhu & Wu, 2018). However, multiple theories suggest a negative correlation including that the early formation of a cold Jupiter may prevent the nuclei of smaller planets from migrating inward (e.g., Izidoro et al., 2015), or that cold Jupiters may reduce the flux of material required to form close-in planets larger than Earth (e.g., Lambrechts et al., 2019). Recently Zhu (2023) looked at the metallicity dimension of the super Earth vs cold Jupiter correlation and found that there is a positive correlation between the two around metal-rich host stars; however, a correlation is unclear for metal-poor host stars due to limited sample size.
@@
Here we analyze the unique WASP-132 planetary system. WASP-132 b was first discovered by Hellier et al. (2017) who measured the planet to have a 7.1 day period, 0.41 0.03 mass, and 0.87 0.03 radius using 23,300 WASP-South (Pollacco et al., 2006) observations from 2006 May - 2012 June, 36 1.2-m Euler/CORALIE (Queloz et al., 2001) RVs from 2014 March - 2016 March, and TRAPPIST (Jehin et al., 2011) photometry data on 2014 May 05. Hord et al. (2022) then later announced the discovery and validation of an inner 1.01 day 1.85 planet using data from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015), but did not characterize the mass. Here we use new RV measurements to characterize the mass of the previously discovered super-Earth WASP-132 c (Hord et al., 2022). We also announce the discovery of a long-period massive giant planet and update all bulk measurements of the system. In Section 2 we present the observations used in this work. In Section 3 we describe our analysis and results. We discuss our results in Section 4 and finally give our conclusions Section 5.
@@ 2.1.1 TESS
As detailed in Hord et al. (2022), WASP-132 (TOI-822, TIC 127530399) was observed by TESS in Sector 11 from UT 2019 April 23 to May 20 (23.96 days) in CCD 2 of Camera 1 and in Sector 38 from UT 2021 April 29 to May 26 (26.34 days) in CCD 1 of Camera 1. Data for WASP-132 were collected at 2-minute cadence in Sectors 11 and 38 and at 20-second cadence in Sector 38. The data were processed by the TESS Science Processing Operation Center (SPOC) pipeline (Jenkins et al., 2016) and were searched by the Transiting Planet Search module (TPS; Jenkins, 2002; Jenkins et al., 2010). The TPS recovered WASP-132 b and WASP-132 c with a period of 1.01153 days with a signal-to-noise ratio (S/N) of 10.6 in the combined data from the two sectors.
Hord et al. (2022) searched and analyzed the TESS light curves using the Presearch Data Conditioning Simple Aperture Photometry (PDC_SAP) TESS light curves generated by the TESS SPOC pipeline (Smith et al., 2012; Stumpe et al., 2012, 2014) at the 2-minute cadence for TESS Sectors 11 and 38. Using the transit least-squares (TLS) search algorithm (Hippke & Heller, 2019), Hord et al. (2022) recovered the hot Jupiter WASP-132 b signal as well as the WASP-132 c signal with a period (1.0119 0.0032 days), depth, and mid-transit time consistent with the values reported by the SPOC pipeline. Hord et al. (2022) also further established the planetary nature of the WASP-132 c signal by using the validation tools vespa (Morton, 2012, 2015) and TRICERATOPS (Giacalone & Dressing, 2020; Giacalone et al., 2021) on the signal and found false-positive probabilities of 9.02 10-5 and 0.0107, respectively.
WASP-132 was recently observed again by TESS in Sector 65 from UT 2023 May 4 to 2023 May 29. Data for WASP-132 were collected at 2-minute cadence in Sector 65. We run an independent analysis of the TESS data, now including Sector 65 and use the PDC_SAP TESS light curves generated by the TESS SPOC pipeline at 2-minute cadence for all three sectors. The full TESS light curves are displayed in Figure 14.
@@ 2.1.2 CHEOPS
We observed three transits of WASP-132 c with the European Space Agency CHaracterizing ExOPlanets Satellite (CHEOPS; Broeg et al., 2013; Benz et al., 2021). CHEOPS is a space telescope with a 32 cm aperture for the purpose of precision follow-up of known planetary systems. We observed transits of WASP-132 c with CHEOPS on 19 May 2022, 4 June 2022, and 11 June 2022 (Program ID: PR430010, PI: B. Akinsanmi) with an exposure time of 60 seconds. The CHEOPS photometry was processed using the dedicated data reduction pipeline (version 13.1; Hoyer et al., 2020). To identify and correct for systematics affecting CHEOPS light curves, we decorrelate the data against the temporal evolution of the telescope roll angle and the flux of background stars inside the photometric aperture (see e.g., Bonfanti et al., 2021). We discuss the removal of systematics in our CHEOPS data in Section 3.3. The CHEOPS light curves are displayed in Figure 15. A table of the reduced CHEOPS photometric observations is available in a machine-readable format at the CDS.
@@ 2.2.1 CORALIE
Observations of WASP-132 with CORALIE began in March 2014 and we have 22 observations with the first upgrade (the ”COR07” setup) and 51 observations after the November 2014 upgrade with the ”COR14” setup. We treat these two data sets as separate instruments to account for any offsets in the measurements. We reduced the spectra with the standard calibration reduction pipeline and computed RVs by cross-correlating with a binary G2 mask (Pepe et al., 2002). We obtained typical RV uncertainties of 21 m s-1 for our CORAIE RVs. The CORALIE RVs are displayed in Figure 1. The CORALIE RVs are presented in tables A.1 and A.2 and are available in a machine-readable format at the CDS.
@@ 2.2.2 HARPS
We obtained 48 HARPS observations of WASP-132 from 17 January 2022 to 1 April 2022 as part of the NOMADS program (PI Armstrong, 108.21YY.001). The HARPS raw data were reduced according to the standard HARPS data reduction software pipeline111ls.eso.org/sci/facilities/lasilla/instruments/harps/doc/index.html with a spectral-type K5 binary mask. We obtained typical RV uncertainties of 2.2 m s-1 for our HARPS RVs. Bisector span and full width at half-maximum (FWHM) of the cross-correlation function (CCF) and other line-profile diagnostics were computed as well. The HARPS spectra were also used to derive spectral parameters for WASP-132, as detailed in Section 3.1. The HARPS RVs are displayed in Figure 1. Figure 1 also displays a two-planet Keplerian model (see Section 3.2.2) for both the CORALIE and HARPS RVs. The HARPS RVs are presented in table A.3 and are available in a machine-readable format at the CDS.
@@ 3.1 Host Star Parameters
Hord et al. (2022) compared several methods to determine the host star parameters of WASP-132 including those reported by Hellier et al. (2017), the TICv8.2 values (Stassun et al., 2018, 2019), an isochrone-based analysis using isoclassify (Huber et al., 2017; Berger et al., 2020), as well as two different broadband Spectral Energy Distribution (SED) plus parallax analyses. Hord et al. (2022) found these methods to be consistent and adopted the isochrone analysis as their final parmameters which used spectroscopic and metallicity from Hellier et al. (2017), Gaia Data Release 2 (DR2; Gaia Collaboration et al., 2016a; Bailer-Jones et al., 2018; Gaia Collaboration et al., 2018a) parallax and coordinates, the Two Micron All-Sky Survey (2MASS; Skrutskie et al., 2006) KS magnitude, and a photometric extinction estimated from Bovy et al. (2016) as inputs.
@@
The spectroscopic and [Fe/H] from Hellier et al. (2017) that Hord et al. (2022) used as inputs for their final model were determined using only 36 CORALIE spectra. We now have 48 higher resolution HARPS spectra and can derive higher precision spectral parameters using a combined HARPS spectrum. The stellar atmospheric parameters (, , microturbulence and [Fe/H]) were derived using the methodology described in Sousa (2014); Santos et al. (2013). We first measured the equivalent widths (EWs) of 224 FeI and 35 FeII lines using the ARES v2 code222The last version of ARES code (ARES v2) can be downloaded at http://www.astro.up.pt/sousasag/ares (Sousa et al., 2015). Then we used these EWs together with a grid of Kurucz model atmospheres (Kurucz, 1993) and the radiative transfer code MOOG (Sneden, 1973) to determine the parameters under assumption of ionization and excitation equilibrium. The abundances of Mg and Si were also derived using the same tools and models as detailed in (e.g. Adibekyan et al., 2012, 2015). Although the EWs of the spectral lines were automatically measured with ARES, we performed careful visual inspection of the EWs measurements as only three lines are available for the Mg abundance determination. Using the combined HARPS spectrum we find a = 468699, [Fe/H] = +0.150.05, and = 4.550.29. We find the star to be a K4V spectral type from its value using the updated table from Pecaut & Mamajek (2013) 333https://www.pas.rochester.edu/~emamajek/EEM_dwarf_UBVIJHK_colors_Teff.txt.
@@
We derived a rotational projected velocity of v = 3.30.6 km s-1 by performing spectral synthesis with MOOG on 36 iron isolated lines and by fixing the stellar parameters, macroturbulent velocity, and limb-darkening coefficient (Costa Silva et al., 2020). The linear limb-darkening coefficient (0.7) was determined using ExoCTK package (Bourque et al., 2021) considering the stellar parameters. We note that a v = 3.30.6 km s-1 yields a Prot/sini = 11.352.09 days (with the value derived below); however, this is not in agreement with the v = 0.90.8 km s-1 Hellier et al. (2017) derived with CORALIE spectra nor with the 333 day signal from a possible rotational modulation in the WASP data. Given that WASP-132 is a cooler star the Doyle et al. (2014) calibrations are not valid, a macroturbulent velocity of 2 km s-1 was assumed and the v value can vary with this assumption.
@@
To derive our final stellar parameters we fit the SED of each star, using the MESA Isochrones and Stellar Tracks (MIST) (Dotter, 2016; Choi et al., 2016) via the IDL suite EXOFASTv2 (Eastman et al., 2019). The stellar parameters are simultaneously constrained by the SED and the MIST isochrones with this method as the SED primarily constrains the stellar radius R⋆ and effective temperature , while a penalty for straying from the MIST evolutionary tracks ensures that the resulting star is physical in nature (see Eastman et al. (2019), for more details on the method).
@@
We use the , , and [Fe/H] values along with their uncertainties from the HARPS spectral analysis as priors for our stellar model and the extinction () is limited to the maximum line-of-sight extinction from the Galactic dust maps of Schlafly & Finkbeiner (2011). We put a Gaussian prior for parallax from the value and uncertainty in DR3 (Gaia Collaboration et al., 2021). We note the DR3 parallax (8.09240) was corrected by subtracting 0.02705 mas according to the Lindegren et al. (2021) prescription. For the SED fit we use photometry from APASS DR9 (Henden et al., 2016); DR3 G, BP, and RP (Gaia Collaboration et al., 2021); 2MASS J, H, and KS (Skrutskie et al., 2006); and ALL-WISE W1, W2, and W3 (Wright et al., 2010), which are presented in Table 1.
@@
We present our final stellar parameters in Table 1. The and [Fe/H] values are from the HARPS spectral analysis, while the , , stellar mass, age, density, and luminosity are outputs of the EXOFASTv2 fit. Overall, we find our stellar parameters including mass and radius to be close in value and within uncertainties of those presented by Hord et al. (2022).

### v2

# Discovery of a cold giant planet and mass measurement of a hot super-Earth in the multi-planetary system WASP-132

@@
Discovery of a cold giant planet and mass measurement of a hot super-Earth in the multi-planetary system WASP-132
Hot Jupiters generally do not have nearby planet companions, as they may have cleared out other planets during their inward migration from more distant orbits. This gives evidence that hot Jupiters more often migrate inward via high-eccentricity migration due to dynamical interactions between planets rather than more dynamically cool migration mechanisms through the protoplanetary disk. Here we further refine the unique system of WASP-132 by characterizing the mass of the recently validated 1.0-day period super-Earth WASP-132 c (TOI-822.02), interior to the 7.1-day period hot Jupiter WASP-132 b. Additionally, we announce the discovery of a giant planet at a 5-year period (2.7 AU). We also detected a long-term trend in the radial velocity data indicative of another outer companion. Using over nine years of CORALIE radial velocities (RVs) and over two months of highly sampled HARPS RVs, we determined the masses of the planets from smallest to largest orbital period to be Mc = , Mb = , and M = , respectively. Using TESS and CHEOPS photometry data, we measured the radii of the two inner transiting planets to be Rc = and Rb = . We find a bulk density of = g cm-3 for WASP-132 c, which is slightly above the Earth-like composition line on the mass-radius diagram. WASP-132 is a unique multi-planetary system in that both an inner rocky planet and an outer giant planet are in a system with a hot Jupiter. This suggests it migrated via a rarer dynamically cool mechanism and helps to further our understanding of how hot Jupiter systems form and evolve.
@@
Some previous studies found that most hot Jupiters do not have nearby planet companions. This was based on non-detections of additional planets in radial velocity (RV) data (e.g., Wright et al., 2009) and photometry data by looking for both additional transit signals (e.g., Steffen et al., 2012; Huang et al., 2016; Hord et al., 2021) and transit-timing variations (TTVs)(e.g., Steffen et al., 2012; Wang et al., 2021; Ivshina & Winn, 2022). The lack of detected planets may be a result of highly eccentric giant planets clearing out any low-mass inner planets in the system during their migration inward (Mustill et al., 2015). These results suggest that high-eccentricity migration due to dynamical interactions between planets (e.g., Rasio & Ford, 1996; Weidenschilling & Marzari, 1996; Lin & Ida, 1997) may be a more common migration mechanism for gas giant planets compared to disk-driven migration (e.g., Goldreich & Tremaine, 1980; Ward, 1997; Baruteau et al., 2014).
@@
However, Wu et al. (2023) recently searched for TTVs across the full four-year Kepler (Borucki et al., 2010) dataset and found that at least 126% of hot Jupiters have a nearby companion. Continued surveys and recent advances in the precision of both photometry and RV observations have allowed the detection of some of these companions, including WASP-47 (Becker et al., 2015), Kepler-730 (Zhu et al., 2018; Cañas et al., 2019), TOI-1130 (Huang et al., 2020), WASP-148 (Hébrard et al., 2020), WASP-132 (Hord et al., 2022), and TOI-2000 (Sha et al., 2023). These companions provide strong evidence that these hot Jupiters migrated via quiescent mechanisms that are dynamically cool and allow nearby planetary companions to remain in the system.
@@
Giant long-period planets have also been detected in systems with small close-in planets (e.g., Santos et al., 2016). Several studies have worked to determine the occurrence of long-period ( 1 year) giant (Mp 0.3 MJup/95 M⊕; Rp 0.45 RJup/5 R⊕) planets, also known as cold Jupiters, in systems that contain close-in (100 days) small (Rp 4 R⊕; Mp 30 M⊕) planets. These studies have varying results with some suggesting a positive correlation between close-in small planets and cold Jupiters (e.g., Zhu & Wu, 2018; Bryan et al., 2019; Herman et al., 2019; Rosenthal et al., 2022), while others find a negative correlation (e.g., Barbato et al., 2018; Bonomo et al., 2023). A positive correlation suggests that these two planet populations do not directly compete for solid material (e.g., Zhu & Wu, 2018). However, multiple theories suggest a negative correlation including that the early formation of a cold Jupiter may prevent the nuclei of smaller planets from migrating inward (e.g., Izidoro et al., 2015), or that cold Jupiters may reduce the flux of material required to form close-in planets larger than Earth (e.g., Lambrechts et al., 2019). Recently, Zhu (2024) looked at the metallicity dimension of the super-Earth versus cold Jupiter correlation and found that there is a positive correlation between the two around metal-rich host stars; however, a correlation is unclear for metal-poor host stars due to a limited sample size.
@@
Here we analyze the unique WASP-132 planetary system. WASP-132 b was first discovered by Hellier et al. (2017) who measured the planet to have a 7.1 day period, 0.41 0.03 mass, and 0.87 0.03 radius using 23,300 WASP-South (Pollacco et al., 2006) observations from 2006 May to 2012 June, 36 1.2-m Euler/CORALIE (Queloz et al., 2001) RVs from 2014 March to 2016 March, and TRAPPIST (Jehin et al., 2011) photometry data on 2014 May 05. Hord et al. (2022) then later announced the discovery and validation of an inner 1.01 day 1.85 planet using data from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015), but did not characterize the mass. Here we use new RV measurements to characterize the mass of the previously discovered super-Earth WASP-132 c (Hord et al., 2022). We also announce the discovery of a long-period massive giant planet, and we updated all bulk measurements of the system. In Sect. 2 we present the observations used in this work. In Sect. 3 we describe our analysis and results. We discuss our results in Sect. 4 and finally give our conclusions Sect. 5.
@@ 2.1.1 TESS
As detailed in Hord et al. (2022), WASP-132 (TOI-822, TIC 127530399) was observed by TESS in Sector 11 from UT 2019 April 23 to May 20 (23.96 days) in CCD 2 of Camera 1 and in Sector 38 from UT 2021 April 29 to May 26 (26.34 days) in CCD 1 of Camera 1. Data for WASP-132 were collected at 2-minute cadence in Sectors 11 and 38 and at 20-second cadence in Sector 38. The data were processed by the TESS Science Processing Operation Center (SPOC) pipeline (Jenkins et al., 2016) and were searched by the Transiting Planet Search module (TPS; Jenkins, 2002; Jenkins et al., 2010). The TPS recovered WASP-132 c with a period of 1.01153 days with a signal-to-noise ratio (S/N) of 10.6 in the combined data from the two sectors.
Hord et al. (2022) searched and analyzed the TESS time series using the Presearch Data Conditioning Simple Aperture Photometry (PDC_SAP) TESS light curves generated by the TESS SPOC pipeline (Smith et al., 2012; Stumpe et al., 2012, 2014) at the 2-minute cadence for TESS Sectors 11 and 38. Using the transit least-squares (TLS) search algorithm (Hippke & Heller, 2019), Hord et al. (2022) recovered the hot Jupiter WASP-132 b signal as well as the WASP-132 c signal with a period (1.0119 0.0032 days), depth, and mid-transit time consistent with the values reported by the SPOC pipeline. Hord et al. (2022) also further established the planetary nature of the WASP-132 c signal by using the validation tools vespa (Morton, 2012, 2015) and TRICERATOPS (Giacalone & Dressing, 2020; Giacalone et al., 2021) on the signal and found false-positive probabilities of 9.02 10-5 and 0.0107, respectively.
WASP-132 was recently observed again by TESS in Sector 65 from UT 2023 May 4 to 2023 May 29. Data for WASP-132 were collected at 2-minute cadence in Sector 65. We run an independent analysis of the TESS data, now including Sector 65 and use the PDC_SAP TESS light curves generated by the TESS SPOC pipeline at 2-minute cadence for all three sectors. The full TESS light curves are displayed in the Supplementary Material published on Zenodo.
@@ 2.1.2 CHEOPS
We observed three transits of WASP-132 c with the European Space Agency CHaracterizing ExOPlanets Satellite (CHEOPS; Broeg et al., 2013; Benz et al., 2021). CHEOPS is a space telescope with a 32 cm aperture for the purpose of precision follow-up of known planetary systems. We observed transits of WASP-132 c with CHEOPS on 19 May 2022, 4 June 2022, and 11 June 2022 (Program ID: PR430010, PI: B. Akinsanmi) with an exposure time of 60 seconds. The CHEOPS photometry was processed using the dedicated data reduction pipeline (version 13.1; Hoyer et al., 2020). To identify and correct for systematics affecting CHEOPS light curves, we decorrelate the data against the temporal evolution of the telescope roll angle and the flux of background stars inside the photometric aperture (see e.g., Bonfanti et al., 2021). We discuss the removal of systematics in our CHEOPS data in Sect. 3.4. Using only CHEOPS data we obtain a transit depth of 588 105 ppm (5.6 detection) for WASP-132 c. The CHEOPS light curves are displayed in the Supplementary Material published on Zenodo.
@@ 2.2.1 CORALIE
Observations of WASP-132 with CORALIE began in March 2014 and we have 22 observations with the first upgrade (the ”COR07” setup) and 51 observations after the November 2014 upgrade with the ”COR14” setup. We treat these two datasets as separate instruments to account for any offsets in the measurements. We reduced the spectra with the standard calibration reduction pipeline and computed RVs by cross-correlating with a binary G2 mask (Pepe et al., 2002). We obtained median RV uncertainties of 21 m s-1 for our CORALIE RVs. We computed stellar activity indicator data from the CORALIE spectra, including: the cross-correlation function (CCF; Baranne et al., 1996) Full-Width at Half-Maximum (FWHM); CCF Bisector velocity span (BIS; Queloz et al., 2001); CCF Contrast, which is the ”depth” of the normalized CCF or height of the Gaussian; and the chromospheric indices measured in H (6562.81 Å) (Boisse et al., 2009), Na (Gomes da Silva et al., 2011), and Ca ii H&K lines (3968.47 and 3933.66 ) (Boisse et al., 2009). The CORALIE RVs are displayed in Fig. 1.
@@ 2.2.2 HARPS
We obtained 48 HARPS observations of WASP-132 from 17 January 2022 to 1 April 2022 as part of the NOMADS program (PI Armstrong, 108.21YY.001). The HARPS raw data were reduced according to the standard HARPS data reduction software pipeline111ls.eso.org/sci/facilities/lasilla/instruments/harps/doc/index.html with a spectral-type K5 binary mask. We obtained median RV uncertainties of 2.2 m s-1 for our HARPS RVs. We computed stellar activity indicators for the HARPS data including CCF FWHM, CCF BIS, contrast of the CCF, H indices, Na indices, and Ca ii indices, and the Mt Wilson ”S index” SMW (Wilson, 1978). The HARPS spectra were also used to derive spectral parameters for WASP-132, as detailed in Sect. 3.1. The HARPS RVs are displayed in Fig. 1. Figure 1 also displays a two-planet Keplerian model plus a linear trend (see Sect. 3.2.2) for both the CORALIE and HARPS RVs.
@@ 3.1 Host star parameters
Hord et al. (2022) compared several methods to determine the host star parameters of WASP-132 including those reported by Hellier et al. (2017), the TICv8.2 values (Stassun et al., 2018, 2019), an isochrone-based analysis using isoclassify (Huber et al., 2017; Berger et al., 2020), as well as two different broadband Spectral Energy Distribution (SED) plus parallax analyses. Hord et al. (2022) found these methods to be consistent and adopted the isochrone analysis as their final parmameters which used spectroscopic and metallicity from Hellier et al. (2017), Gaia Data Release 2 (DR2; Gaia Collaboration et al., 2016; Bailer-Jones et al., 2018; Gaia Collaboration et al., 2018) parallax and coordinates, the Two Micron All-Sky Survey (2MASS; Skrutskie et al., 2006) KS magnitude, and a photometric extinction estimated from Bovy et al. (2016) as inputs.
@@
The spectroscopic and [Fe/H] from Hellier et al. (2017) that Hord et al. (2022) used as inputs for their final model were determined using only 36 CORALIE spectra. We now have 48 higher resolution HARPS spectra and can derive higher precision spectral parameters using a combined HARPS spectrum. The stellar atmospheric parameters (, , microturbulence and [Fe/H]) were derived using the methodology described in Sousa (2014); Santos et al. (2013). We first measured the equivalent widths (EWs) of 224 Fe I and 35 Fe II lines using the ARES v2 code222The last version of ARES code (ARES v2) can be downloaded at http://www.astro.up.pt/sousasag/ares (Sousa et al., 2015). Then we used these EWs together with a grid of Kurucz model atmospheres (Kurucz, 1993) and the radiative transfer code MOOG (Sneden, 1973) to determine the parameters under assumption of ionization and excitation equilibrium. The abundances of Mg and Si were also derived using the same tools and models as detailed in (e.g. Adibekyan et al., 2012, 2015). Although the EWs of the spectral lines were automatically measured with ARES, we performed careful visual inspection of the EWs measurements as only three lines are available for the Mg abundance determination. Using the combined HARPS spectrum we find a = 468699 K, [Fe/H] = +0.150.05, and = 4.550.29 (cgs). We find the star to be a K4V spectral type from its value using the updated table from Pecaut & Mamajek (2013) 333https://www.pas.rochester.edu/~emamajek/EEM_dwarf_UBVIJHK_colors_Teff.txt.
@@
We derived the rotational projected velocity by performing spectral synthesis with MOOG on 36 iron isolated lines, using Gaussian priors for the spectroscopic parameters, and fixing the macroturbulent velocity and limb-darkening coefficient (Costa Silva et al., 2020). The linear limb-darkening coefficient (0.7) was determined using ExoCTK package (Bourque et al., 2021) considering the stellar parameters. Given that WASP-132 is a cooler star the Doyle et al. (2014) calibrations are not valid. We assumed a macroturbulent velocity of 2 km s-1, but the v value can vary with this assumption. With this method we find a rotational projected velocity of v = 3.30.6 km s-1. We note that a v = 3.30.6 km s-1 yields a Prot/sini = 11.352.09 days (with the value derived below); however, this is not in agreement with the v = 0.90.8 km s-1. Hellier et al. (2017) derived with CORALIE spectra nor with the 333 day signal from a possible rotational modulation in the WASP data. We tested larger values for macroturbulent velocities and found v values of 2.9, 2.4, and 1.5 km s-1 for macroturblent velocities of 3, 3.5, and 4 km s-1, respectively. Given the large variance of v with the unknown macroturbulent velocity the v = 3.30.6 km s-1 can only be taken as an estimate.
@@
To derive our final stellar parameters, including the age of the star, we fit the SED of the star using MESA Isochrones and Stellar Tracks (MIST) (Dotter, 2016; Choi et al., 2016) via the IDL suite EXOFASTv2 (Eastman et al., 2019). The stellar parameters are simultaneously constrained by the SED and the MIST isochrones with this method as the SED primarily constrains the stellar radius and effective temperature , while a penalty for straying from the MIST evolutionary tracks ensures that the resulting star is physical in nature (see Eastman et al. (2019), for more details on the method).
@@
We use the , , and [Fe/H] values along with their uncertainties from the HARPS spectral analysis as priors for our stellar model and the extinction () is limited to the maximum line-of-sight extinction from the Galactic dust maps of Schlafly & Finkbeiner (2011). We put a Gaussian prior for parallax from the value and uncertainty in DR3 (Gaia Collaboration et al., 2021). We note the DR3 parallax (8.09240 mas) was corrected by subtracting 0.02705 mas according to the Lindegren et al. (2021) prescription. For the SED fit we use photometry from APASS DR9 (Henden et al., 2016); DR3 G, BP, and RP (Gaia Collaboration et al., 2021); 2MASS J, H, and KS (Skrutskie et al., 2006); and ALL-WISE W1, W2, and W3 (Wright et al., 2010), which are presented in Table 1.
@@
We present our final stellar parameters in Table 1. The and [Fe/H] values are from the HARPS spectral analysis, while the , , stellar mass, age, density, luminosity, and distance are outputs of the EXOFASTv2 fit. Overall, we find our stellar parameters including mass and radius to be close in value and within uncertainties of those presented by Hord et al. (2022), as well as a close agreement between the HARPS spectroscopic derived and that derived from EXOFASTv2.